# Understanding How Psychological Distance Influences User Preferences in Conversational Versus Web Search

### v1

@@ Abstract.
Conversational search offers an easier and faster alternative to conventional web search, while having downsides like lack of source verification. Research has examined performance disparities between these two systems in different settings. However, little work has considered the effects of variations within a given search task. We hypothesize that psychological distance - one’s perceived closeness to a target event - affects information needs in search tasks, and investigate the corresponding effects on user preferences between web and conversational search systems. We find that with greater psychological distances, users perceive conversational search as more credible, useful, enjoyable, and easy to use, and demonstrate increased preference for this system. We reveal qualitative reasons for these differences and provide design implications for search system designers.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCIccs: Information systems Search interfacesccs: Computing methodologies Artificial intelligenceccs: Information systems Web searching and information discovery
@@ 1. Introduction
However, a crucial overlooked factor is the inherent variation in individual tasks. Notably, tasks can vary significantly in terms of the perceived psychological distance to the user: for instance, whether they are seeking information for themselves or someone else, or whether they are searching about an event more or less likely to happen. The well-established Construal Level Theory (CLT) describes the relation between psychological distance and the level of abstraction of thought (Trope and Liberman, 2010). The psychological distance of a task (e.g. planning a trip for next week versus next year) influences whether people think about specific details (e.g. public transport options) or general ideas (e.g. cities to visit) (Trope and Liberman, 2010). Such variations significantly impact how users process information and make decisions (Lermer et al., 2015; Trope and Liberman, 2010; Wakslak and Trope, 2009; Fiedler, 2007); we propose that this effect extends to information needs in IR tasks, and thus user perceptions and preferences of the IR systems that fulfil them. Psychological distance is a common factor in everyday life, being assessed automatically without conscious effort (Liberman and Trope, 2014), and has fundamental and widespread importance for cognitive and behavioral decision making (Liberman and Trope, 2014; Maglio et al., 2013). Therefore, the effect of psychological distance of IR tasks on preferences for IR systems presents a key research gap.
@@
To investigate this, we conducted a 2x4 between-subjects study, with 128 participants divided into eight groups. Each participant was assigned to one of four psychological distance dimensions posited by CLT — spatial (where an event takes place), temporal (when an event occurs), social (who is involved in the event), or hypothetical (how likely the event is to happen) — and one of two psychological distance levels (near or far). Participants completed an IR task, manipulated according to their assigned dimension and distance, using both web and conversational search. They were then asked to rate the perceived credibility, perceived usefulness, enjoyability, and ease of use of each tool, and finally their preference between tools.
@@
Our findings indicated that participants in the far psychological distance condition exhibited a significantly higher level of preference for the LLM tool over traditional web search, reporting greater perceived credibility, perceived usefulness, enjoyment, and ease of use. Additionally, participants’ search behaviors varied depending on psychological distance. When distance was perceived as greater, participants more effectively utilized the strengths of the LLM system to explore broad ideas, such as cities, cuisines, and accommodations. In contrast, nearer distances prompted participants to focus on concrete steps for planning and executing their trips. These findings suggest that the choice of IR system is closely tied to the user’s situational context, and offer important design implications for designers of IR systems.
@@
    It advances our understanding of the differences in user perceptions and preferences when interacting with information through traditional search engines and LLM-powered conversational search, contributing to the literature on Human-Information Interaction. Additionally, we explore how these preferences can inform tool recommendations based on psychological distance for information-seeking.
@@ 2. Related Work
2.1. Information Seeking
@@ 2.2. Construal Level Theory and Psychological Distance
To explore how specific tasks influence user preferences for web or conversational search, we turn to CLT, which provides a theoretical framework for manipulating task situation through psychological distance. Psychological distance is a common and critical factor in information seeking tasks, such as planning for near or distant events, searching information for oneself or someone else. These variations in psychological distance can significantly impact how users process information and make decisions (Lermer et al., 2015; Trope and Liberman, 2010; Wakslak and Trope, 2009; Fiedler, 2007). CLT not only provides ways to manipulate psychological distance of tasks, but also provides a theoretical basis for the potential impact on user preferences towards different IR systems. When users make choices or set preferences for different objects, they rely on their understanding of these objects (Trope and Liberman, 2010). In CLT, increasing psychological distance enhances higher-level understanding, which in turn influences user evaluations and preferences (Trope and Liberman, 2010). CLT posits that as the psychological distance of a mental phenomenon increases, the construal level of thought processes shifts from low-level concrete constructions to high-level abstract constructions (Trope and Liberman, 2010). We believe that in the context of information seeking, the perceived psychological distance of an information seeking task presented to a user affects their construal level of thinking. Task psychological distances may be smaller (e.g. finding information for oneself) or greater (e.g. finding information for someone else). CLT also discusses how psychological distance affects task-related evaluations (Trope and Liberman, 2010). As psychological distance increases, users tend to seek more supportive viewpoints than opposing ones (Trope and Liberman, 2010). This is because higher-level thinking prioritizes the pros of a situation over its cons (Trope and Liberman, 2010). Herzog et al. (Herzog et al., 2007) found that as psychological distance increases, it becomes easier for users to generate supportive viewpoints and harder to generate opposing ones. Thus, we propose that by varying the psychological distances of tasks, users will perceive the nature of their task differently via different construal levels(Trope and Liberman, 2012), and hence tend toward either web search or conversational search.
@@
CLT proposes four main dimensions of psychological distance: spatial, temporal, social, and hypothetical (Trope and Liberman, 2010). Spatial distance is where events are going to happen; temporal distance concerns when events are going to happen; social distance is how similar the people involved in an event are to oneself; and hypothetical distance is the likelihood of an event happening (Trope and Liberman, 2010). Each dimension has some unique characteristics, but all four influence the mental models of a user through their construal level (Liberman and Trope, 2014). Low-level constructions are unstructured and detail-rich, encompassing primary, secondary, and incidental characteristics of an event (Trope and Liberman, 2010). In contrast, high-level constructions are more abstract and less specific, extracting essential points from available information while omitting details (Trope and Liberman, 2010). In comparing web search and conversational search, we propose that web search aligns more with the need for low-level constructions, while LLM-powered conversational search meets higher-level construction needs. Therefore, our study is based on the theory that the psychological distance associated with information seeking tasks influences user information need through different levels of construal. The information seeking activities users engage in to satisfy their current information needs, in turn, affect their perceptions and preferences for these two technologies.
2.3. Conversational Search as a New Means of IR
IR is the task of searching for relevant information from an online repository to fulfil a user need. Conversational search, a specialized branch of IR systems, aims to help users obtain information that meets their needs through flexible and natural dialogues, based on their queries and relevant feedback (Kiesel et al., 2021; Jannach et al., 2021; Liao et al., 2020; Radlinski and Craswell, 2017; Croft and Thompson, 1987). Modern LLMs provide a revolutionary step forward in empowering such systems and their advantages, due to their outstanding performance in understanding and generating human-like text (Gao et al., 2020).
Conversational search offers a number of innate advantages over traditional web search. Unlike web search queries that require specific terms and keywords, conversational search enables users to express their needs through natural dialogue, reducing the effort needed to clarify their particular aims (Radlinski and Craswell, 2017). Conversational systems also significantly reduce cognitive load (Kaushik, 2019) by providing summaries instead of the raw content of traditional search engines (Kaushik and Jones, 2023). For search tasks that involve multiple query aspects, such as making travel plans that include accommodation, food, and attractions, conversational search can be preferable in allowing users to assess the overall utility of such multi-item compositions (Radlinski and Craswell, 2017). Furthermore, the natural conversational abilities and human-like qualities of LLMs provide greater social interaction, which itself is valuable in enhancing the overall search experience (Følstad and Brandtzaeg, 2020).
However, LLM-powered conversational search also suffers from a few disadvantages. LLMs, as natural language generation models (Brown et al., 2020), inevitably experience hallucination (Maynez et al., 2020), generating information that contradicts or is not supported by their source content (Ji et al., 2023). Moreover, LLMs often do not provide specific sources for the summaries they generate (Bender et al., 2021). Finally, because their training data are based on past internet information with a specific cut-off point, (Achiam et al., 2023), the information provided by LLMs can be outdated, and they may be unable to answer queries that require the latest data or information (Liu et al., 2023). This therefore raises the question of when LLM may or may not be superior to traditional web search.
@@ 2.4. Psychological Distance in Information Seeking Tasks
Previous research has highlighted the potential advantages of conversational search in terms of ease of use (reducing user effort), usefulness (especially for complex tasks), and enjoyability, as well as potential issues with credibility. However, our exploration of user preferences for conversational versus web search considers not only the impact of the interactive aspect, but also the nature of the information seeking tasks. As mentioned in section 2.1, the nature of information seeking tasks influence users’ choices of information sources, which should affect their preferences for different IR systems.
Previous studies have investigated the relative efficacy of conversational search and web search across various types of information seeking tasks, such as healthcare and tourism. Changes in task type can alter user information needs, with conversational search and web search differing in their ability to meet these needs. Specifically, in the healthcare context, studies have demonstrated that generative pre-trained transformers (GPTs) surpass traditional web searches like Google in diagnosing prevalent illnesses and suggesting subsequent treatment steps (Hristidis et al., 2023; Shahsavar et al., 2023; Sandmann et al., 2024). In the realm of tourism, ChatGPT has been shown to significantly enhance the planning of trips. They serve as virtual assistants, offering personalized advice, and contribute to a richer travel experience by integrating diverse information sources (Dwivedi et al., 2024; Zarezadeh et al., 2023; Wong et al., 2023; Zhang et al., 2024). Despite these advancements, these studies also highlight the limitations in conversational search, closely related to the challenges associated with LLMs mentioned in section 2.3, underscoring the importance of traditional web search (Sandmann et al., 2024). Users frequently turn to web search to verify LLM-generated information, indicating that LLMs complement rather than replace web search in information seeking tasks (Yen et al., 2024).
@@
However, even if the type of task remains constant, psychological distance, a common factor in information seeking tasks, can also lead to changes in user information needs. Few studies have explored the impact of psychological distance in information seeking tasks. Moreover, previous studies have rarely focused on user perceptions and preferences for these two IR systems, lacking a comprehensive perspective from task nature to technology preference. Therefore, this study aims to fill this gap by considering the psychological distance in information seeking tasks to deepen our understanding of user preferences for conversational search versus web search in real-world information searching. This not only helps us to understand the motivations behind user choices at specific task psychological distances, but also provides important insights for designing IR systems that better meet user needs. Since CLT provides a theoretical foundation for both manipulating tasks through the critical factor (psychological distances) and understanding their impact on user preferences, we investigate this gap by manipulating the psychological distance of information seeking tasks using CLT. We are particularly interested in exploring:
Web search conventionally provides a list of results ranked according to the relevance of the user’s query, allowing users to access specific information through links (Brin and Page, 1998). However, obtaining a comprehensive overview of a topic is more difficult as users must access and filter through multiple sources and aggregate this information themselves. Conversely, LLM-powered conversational search generates summaries of the information relevant to the user’s current search intent, providing a more comprehensive overview of relevant information. However, these tend to lack specific details (Achiam et al., 2023; Brown et al., 2020; Devlin et al., 2018), and require user to ask further questions or refine their prompts for more detailed inquiries. Further, when users require highly-specific or time-sensitive information, LLM-powered conversational search may not provide accurate answers (e.g. local train ticket prices or current weather conditions).
CLT posits that tasks with greater psychological distances elicit preferences for more abstract, higher-levels overviews of information. Thus, we propose that the perceived ease of use of each IR system depends on the psychological distance of the task and thus the needs of the user. We hypothesize:
CLT suggests that as psychological distance increases, users tend to seek more supportive rather than opposing viewpoints, due to high-level thinking’s tendency to overlook specific disadvantages. In conversational search, recent studies (Sharma et al., 2024) suggest that LLM-powered conversational search creates an echo chamber effect, producing content that validates user views more than web search and reducing the occurences of opposing viewpoints. Content that corroborates user views is perceived as more credible (Nickerson, 1998), specifically in terms of accuracy, authenticity, and believability (Appelman and Sundar, 2016). Therefore, we hypothesize the following.
@@
Extensions of the TAM have identified additional relevant variables beyond perceived usefulness and ease of use. One such factor is perceived enjoyment (Davis et al., 1992) expanded the TAM to include perceived enjoyment, which can also directly enhance user attitudes and perceived usefulness: when users find a technology enjoyable, they also perceive it as more useful, and vice versa. In another extension of the TAM for consumer-generated media that aids travel planning, Ayeh (Ayeh, 2015) introduced information credibility as a prerequisite for perceived usefulness. Applying the TAM and its extensions in conversational search, we illustrate the interactions among user perceptions of these systems in Figure 1. Based on H1 and H2, we propose the following hypotheses:

### v2

@@ Abstract.
Conversational search offers an easier and faster alternative to conventional web search, while having downsides like a lack of source verification. Research has examined performance disparities between these two systems in various settings. However, little work has investigated how changes in the nature of a search task affect user preferences. We investigate how psychological distance - the perceived closeness of one to an event - affects user preferences between conversational and web search. We hypothesise that tasks with different psychological distances elicit different information needs, which in turn affect user preferences between systems. Our study finds that, under fixed condition ordering, greater psychological distances lead users to prefer conversational search, which they perceive as more credible, useful, enjoyable, and easy to use. We reveal qualitative reasons for these differences and provide design implications for search system designers.
@@
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713770isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in HCIccs: Information systems Search interfacesccs: Computing methodologies Artificial intelligenceccs: Information systems Web searching and information discovery
@@ 1. Introduction
However, a crucial overlooked factor is the inherent variation in individual tasks. Tasks can vary significantly in terms of the perceived psychological distance to the user: for instance, whether they are seeking information for themselves or someone else, or whether they are searching about an event more or less likely to happen. The well-established Construal Level Theory (CLT) describes the relation between psychological distance and the level of abstraction of thought (Trope and Liberman, 2010). The psychological distance of a task (e.g. planning a trip for next week versus next year) influences whether people think about specific details (e.g. public transport options) or general ideas (e.g. cities to visit) (Trope and Liberman, 2010). Such variations significantly impact how users process information and make decisions (Lermer et al., 2015; Trope and Liberman, 2010; Wakslak and Trope, 2009; Fiedler, 2007); we propose that this effect extends to information needs in IR tasks, and thus user perceptions and preferences of the IR systems that fulfill them. Psychological distance is a common factor in everyday life, being assessed automatically without conscious effort (Liberman and Trope, 2014), and has fundamental and widespread importance for cognitive and behavioral decision making (Liberman and Trope, 2014; Maglio et al., 2013). Therefore, the effect of psychological distance of IR tasks on preferences for IR systems presents a key research gap.
@@
To investigate this, we conducted a 2x4 between-subjects study, with 128 participants divided into eight groups. Each participant was assigned to one of four psychological distance dimensions posited by CLT — spatial (where an event takes place), temporal (when an event occurs), social (who is involved in the event), or hypothetical (how likely the event is to happen) — and one of two psychological distance levels (near or far). Participants completed an IR task, manipulated according to their assigned dimension and distance, first using web search and then conversational search. They were then asked to rate the ease of use, enjoyment, and perceived credibility of each tool, and finally their preference between tools.
@@
Our findings indicated that participants in the far psychological distance condition exhibited a significantly higher level of preference for the LLM tool over traditional web search, reporting greater ease of use, usefulness, enjoyment, and credibility. Additionally, participants’ search behaviors varied depending on psychological distance. When distance was perceived as greater, participants more effectively utilized the strengths of the LLM system to explore broad ideas, such as cities, cuisines, and accommodations. In contrast, nearer distances prompted participants to focus on concrete steps for planning and executing their trips. These findings suggest that the choice of IR system is closely tied to the user’s situational context, and offers important design implications for designers of IR systems.
@@
    It advances our understanding of the differences in user perceptions and preferences when interacting with information through traditional search engines and LLM-powered conversational search, contributing to the literature on Human-Information Interaction. Additionally, we explore how these preferences can inform tool recommendations based on psychological distance for information seeking.
@@ 2. Related Work
2.1. Conversational Search as a New Means of IR for Information Seeking
@@
IR is the task of searching for relevant information from an online repository to fulfill a user’s need. Conversational search, a specialized branch of IR systems, aims to help users obtain information that meets their needs through flexible and natural dialogues, based on their queries and relevant feedback (Kiesel et al., 2021; Jannach et al., 2021; Liao et al., 2020; Radlinski and Craswell, 2017; Croft and Thompson, 1987). LLMs provide a revolutionary step forward in empowering such systems and their advantages due to their outstanding performance in understanding and generating human-like text (Gao et al., 2020).
Conversational search offers a number of innate advantages over traditional web search. Unlike web search queries that require specific terms and keywords, conversational search enables users to express their needs through natural dialogue, reducing the effort needed to clarify their particular aims (Radlinski and Craswell, 2017) and reducing cognitive load (Kaushik, 2019). For example, (Kaushik and Jones, 2023) found that users perceived less cognitive load and greater usability with a conversational search interface, compared to a standard search interface with the same knowledge base. For search tasks that involve multiple query aspects, such as making travel plans that include accommodation, food, and attractions, conversational search can be preferable in allowing users to assess the overall utility of such multi-item compositions (Radlinski and Craswell, 2017). Furthermore, the natural conversational abilities and human-like qualities of LLMs provide greater social interaction, which itself is valuable in enhancing the overall search experience (Følstad and Brandtzaeg, 2020).
However, LLM-powered conversational search also suffers from a few disadvantages. LLMs, as natural language generation models (Brown et al., 2020), inevitably experience hallucination (Maynez et al., 2020), generating information that contradicts or is not supported by their source content (Ji et al., 2023). Moreover, LLMs often do not provide specific sources for the summaries they generate (Bender et al., 2021). Finally, because their training data are based on past internet information with a specific cut-off point (Achiam et al., 2023), the information provided by LLMs can be outdated, and they may be unable to answer queries that require the latest data or information (Liu et al., 2023). This therefore raises the question of when LLMs may or may not be superior to traditional web search. We address this issue in this study by asking participants to directly compare the two search systems in a realistic use case.
@@ 2.2. Construal Level Theory and Psychological Distance
To explore how specific tasks influence user preferences for web or conversational search, we turn to CLT, which provides a theoretical framework for manipulating task situation through psychological distance. Psychological distance is a common and critical factor in information seeking tasks, such as planning for near or distant events, or searching information for oneself or someone else. These variations in psychological distance can significantly impact how users process information and make decisions (Lermer et al., 2015; Trope and Liberman, 2010; Wakslak and Trope, 2009; Fiedler, 2007). CLT not only provides ways to manipulate the psychological distance of tasks, but also provides a theoretical basis for the potential impact on user preferences towards different IR systems. When users make choices or set preferences for different objects, they rely on their understanding of these objects (Trope and Liberman, 2010). In CLT, increasing psychological distance enhances higher-level understanding, which in turn influences user evaluations and preferences (Trope and Liberman, 2010). CLT posits that as the psychological distance of a mental phenomenon increases, the construal level of thought processes shifts from low-level concrete constructions to high-level abstract constructions (Trope and Liberman, 2010).
CLT suggests that as distance and hence construal levels shift, differing mental representations of a given task affect how people approach that task (Trope and Liberman, 2010) - specifically, they affect the information and features that people find salient and relevant to task completion. For instance, distance was found to influence purchase-making decisions (Trope and Liberman, 2000), with closer distances causing participants to prioritize central features of the product over its peripheral features. Similar effects were found in evaluation of student essays (Liviatan et al., 2008) and advice given for choosing a job (Kray and Gonzalez, 1999). We propose that psychological distance also affects the way users approach information seeking tasks. Since tasks with different psychological distances cause users to focus on obtaining different kinds of information, and web and conversational search fulfill different information needs, varying the psychological distances of tasks could cause differing preferences between the two search systems.
@@
CLT proposes four main dimensions of psychological distance: spatial, temporal, social, and hypothetical (Trope and Liberman, 2010). Spatial distance is where events are going to happen; temporal distance concerns when events are going to happen; social distance is how similar the people involved in an event are to oneself; and hypothetical distance is the likelihood of an event happening (Trope and Liberman, 2010). Each dimension has some unique characteristics, but all four influence the mental models of a user through their construal level (Liberman and Trope, 2014). Low-level constructions are unstructured and detail-rich, encompassing primary, secondary, and incidental characteristics of an event (Trope and Liberman, 2010). In contrast, high-level constructions are more abstract and less specific, extracting essential points from available information while omitting details (Trope and Liberman, 2010). We propose that web search aligns more with the need for low-level constructions, while LLM-powered conversational search meets higher-level construal needs. Therefore, as the psychological distance associated with an information seeking task influences user information needs and information seeking behaviors through different construal levels, this in turn affects their perceptions and preferences for these two technologies. We explore this potential novel difference by presenting participants with tasks of different psychological distances, and comparing the resulting differences in user preference.
@@ 2.3. Psychological Distance in Information Seeking Tasks
Previous research has highlighted the potential advantages of conversational search in terms of ease of use (reducing user effort), usefulness (especially for complex tasks), and enjoyability, as well as potential issues with credibility. For user preferences, the effects of these interactive aspects are certainly important. In addition, the nature of the information seeking tasks is also an aspect worth considering. As mentioned in section 2.1, the nature of information seeking tasks influences users’ choices of information sources, which should affect their preferences for different IR systems. Previous studies have explored the relative efficacy of conversational search and web search across various types of information seeking tasks, particularly in domains such as healthcare and tourism. In the healthcare context, Hristidis et al. (Hristidis et al., 2023) demonstrated that ChatGPT provides more relevant and objective responses to dementia-related queries compared to Google, though it lacks source reliability and timestamped information. Similarly, Shahsavar et al. (Shahsavar et al., 2023) identified that users are increasingly inclined to use ChatGPT for self-diagnosis due to its higher performance expectancy and positive risk-reward appraisals, despite concerns about its suitability for healthcare purposes. Sandmann et al. (Sandmann et al., 2024) further highlighted that GPT-4 outperforms Google in diagnostic accuracy across diverse clinical cases, showcasing the growing potential of LLMs in medical question answering. In the tourism domain, Zarezadeh et al. (Zarezadeh et al., 2023) emphasized the potential of AI tools like ChatGPT to streamline information searches. Wong et al. (Wong et al., 2023) illustrated ChatGPT’s ability to enhance decision-making during different stages of travel by providing personalized and cost-effective recommendations. Zhang et al. (Zhang et al., 2024) extended this by showcasing ChatGPT 4.0’s multimodal capabilities, enabling users to compare destinations and plan activities. Finally, Dwivedi et al. (Dwivedi et al., 2024) outlined the transformative potential of generative AI in the hospitality and tourism industry while identifying challenges related to bias and data privacy. Despite these advancements, these studies also highlight the limitations in conversational search, closely related to the challenges associated with LLMs mentioned in section 2.1, underscoring the importance of traditional web search (Sandmann et al., 2024). Users frequently turn to web search to verify LLM-generated information, indicating that LLMs complement rather than replace web search in information seeking tasks (Yen et al., 2024).
@@
However, even if the type of task remains constant, psychological distance, a common factor in information seeking tasks, can also lead to changes in user information needs. Few studies have explored the impact of psychological distance in information seeking tasks. Moreover, previous studies have rarely focused on user perceptions and preferences for these two IR systems, lacking a comprehensive perspective from task nature to technology preference. Therefore, this study aims to fill this gap by considering the psychological distance in information seeking tasks to deepen our understanding of user preferences for conversational search versus web search in real-world information searching. This not only helps us to understand the motivations behind user choices at specific task psychological distances, but also provides important insights for designing IR systems that better meet user needs.
3. Current Study
@@
As mentioned in section 2.2, CLT provides a theoretical foundation for both manipulating tasks through the critical factor (psychological distances) and understanding their impact on user preferences. We investigate this gap by manipulating the psychological distance of information seeking tasks using CLT. We are particularly interested in exploring:
Web search conventionally provides a list of results ranked according to the relevance of the user’s query, allowing users to access specific information through links (Brin and Page, 1998). However, obtaining a comprehensive overview of a topic is more difficult as users must access and filter through multiple sources and aggregate this information themselves. Conversely, LLM-powered conversational search generates summaries of the information relevant to the user’s current search intent, providing a more comprehensive overview of relevant information. However, these tend to lack specific details (Achiam et al., 2023; Brown et al., 2020; Devlin et al., 2018), and require users to ask further questions or refine their prompts for more detailed inquiries. Further, when users require highly-specific or time-sensitive information, LLM-powered conversational search may not provide accurate answers (e.g. local train ticket prices or current weather conditions).
CLT posits that tasks with greater psychological distances elicit preferences for more abstract, higher-levels overviews of information. Thus, we propose that the perceived ease of use of each IR system depends on the psychological distance of the task and thus the needs of the user. We hypothesise:
CLT suggests that as psychological distance increases, users tend to seek more supportive rather than opposing viewpoints, due to high-level thinking’s tendency to overlook specific disadvantages. In conversational search, recent studies (Sharma et al., 2024) suggest that LLM-powered conversational search creates an echo chamber effect, producing content that validates user views more than web search and reducing the occurrences of opposing viewpoints. Content that corroborates user views is perceived as more credible (Nickerson, 1998), specifically in terms of accuracy, authenticity, and believability (Appelman and Sundar, 2016). Therefore, we hypothesise the following.
Extensions of the TAM have identified additional relevant variables beyond perceived usefulness and ease of use. One such factor is perceived enjoyment. Davis et al. (Davis et al., 1992) expanded the TAM to include perceived enjoyment, which can also directly enhance user attitudes and perceived usefulness: when users find a technology enjoyable, they also perceive it as more useful, and vice versa. In another extension of the TAM for consumer-generated media that aids travel planning, Ayeh (Ayeh, 2015) introduced information credibility as a prerequisite for perceived usefulness. In the context of IR, credibility plays an important role by indirectly influencing user preferences. Users are more likely to adopt and prefer systems that provide reliable information, as credibility directly affects their trust in the system (McKnight and Kacmar, 2007). When users trust the information provided by the system, they are more likely to perceive it as useful (Ayeh, 2015; Ayeh et al., 2013; McKnight and Kacmar, 2007), which is consistent with the core structure of the TAM model (Davis, 1989). Applying the TAM and its extensions in conversational search, we illustrate the interactions among user perceptions of these systems in Figure 1. Based on H1 and H2, we propose the following hypotheses: